OpenStudy (anonymous):

f(x) = 4x + 5 h(x) = √5x - 4 Find: (hf)(4)

OpenStudy (anonymous):

Those are the choices √17 84 5 25

OpenStudy (anonymous):

Sweeeeet I love these.. Hold up

OpenStudy (anonymous):

ok i have a lot of them lol

OpenStudy (anonymous):

If i get this wrong i'm gonna feel bad :(

OpenStudy (anonymous):

This is a hard one

OpenStudy (anonymous):

lol well i gave choices

OpenStudy (anonymous):

If i take my best guess will you be mad at me ???? :)

OpenStudy (anonymous):

lol is it even close to any of the choices?

OpenStudy (espex):

For clarification, is your h(x) equation \[\sqrt{5x-4}\] or \[\sqrt{5x}-4\]

OpenStudy (espex):

Either way, the basic principle here is that you have two functions multiplied together and you substitute '4' in every place you have an 'x'.

OpenStudy (anonymous):

its over the whole thing

OpenStudy (anonymous):

and i tried it and i got 25

OpenStudy (espex):

I do not see how that is possible. Could you please show me how you got it?

OpenStudy (anonymous):

i did f(4)=4(4) +5 16+5 =21 then i did h(4) = sqaure root thing 5(4)-4 then sqaure root 16 = 4 then added 21 and 4

OpenStudy (espex):

You are right all the way up to the last step. You are not being asked to solve (f+h)(x), you are given (f*h)(x)

OpenStudy (anonymous):

so 21 * 4?

OpenStudy (espex):

Yes.

OpenStudy (anonymous):

oh!! so 84 yay i got it

OpenStudy (espex):

Good job. :)

OpenStudy (anonymous):

what about this f(x) = 2x - 7 g(x) = -4x2 - 6x Find: g(f(x)) -8x3 + 16x2 + 42x -8x2 - 12x - 7 -16x2 + 100x - 154 20x2 + 30x

OpenStudy (espex):

As I'm sure you have guessed, this one does not have a value for 'x' so you won't be able to solve to completion. However the same principle applies only this time you are putting the 'f(x)' into 'g' as the 'x' component.

OpenStudy (anonymous):

That was my first guess...

OpenStudy (espex):

What if we made an example and said, f(x)=4, find g(f(x))? What would you do?

OpenStudy (anonymous):

i got -8x3 + 16x2 + 42x

OpenStudy (espex):

How?

OpenStudy (anonymous):

(-4x^2 -6x)(2x-7) -8^3 +28x^2 -12x^2 +42 -8^3 +16x^2 +42

OpenStudy (espex):

You are not multiplying these two expressions.

OpenStudy (espex):

If I gave you g(x)=4x+3, x=2, and asked you to solve this, how would you do it?

OpenStudy (espex):

You would plug the 2 in everywhere you saw an 'x' and work the expression.

OpenStudy (anonymous):

the one i just typed was the answer to this f(x) = 2x - 7 g(x) = -4x2 - 6x Find: g(f(x))

OpenStudy (espex):

That is exactly what you are asked to do here, only instead of '2' you must plug in '2x-7'

OpenStudy (espex):

I would not agree that 8x^3+16x^2+42 was the answer.

OpenStudy (anonymous):

im confused.......

OpenStudy (espex):

Okay, you are given g(x) = -4x2 - 6x, f(x) = 2x - 7, and asked to find g(f(x)). Let's try a mental exercise before we attack this problem. What if instead I gave you g(x) = -4x^2-6x and told you that f(x)=2, then asked you to find g(f(x)), how would you approach solving this?

OpenStudy (anonymous):

g(f(2)=-4(2)^2-6(2)

OpenStudy (espex):

Yes! The right idea, off a little on notation but we can handle that.

OpenStudy (espex):

So now the problem you have, g(x) = -4x2 - 6x, f(x) = 2x - 7, and asked to find g(f(x)), do the same thing only instead of plugging in the '2' you will now use '2x-7'

OpenStudy (anonymous):

I think you need to go a little "slower" on her.

OpenStudy (anonymous):

g(f(2x-7))?!!?!?!?

OpenStudy (espex):

I don't think so, she has the idea and is clearly capable of understanding. We just need to make that connection for her on this and she's nailed it.

OpenStudy (espex):

Exactly, although the proper notation would be g(2x-7) and so you would plug '2x-7' in everywhere there is an 'x' in your 'g' expression.

OpenStudy (anonymous):

Or are you trying to nail it??

OpenStudy (anonymous):

fuknut